### v1

# CountingFruit: Real-Time 3D Fruit Counting with Language-Guided Semantic Gaussian Splatting

Abstract
Accurate fruit counting in real-world agricultural environments is a longstanding challenge due to visual occlusions, semantic ambiguity, and the high computational demands of 3D reconstruction. Existing methods based on neural radiance fields suffer from low inference speed, limited generalization, and lack support for open-set semantic control. This paper presents FruitLangGS, a real-time 3D fruit counting framework that addresses these limitations through spatial reconstruction, semantic embedding, and language-guided instance estimation. FruitLangGS first reconstructs orchard-scale scenes using an adaptive Gaussian splatting pipeline with radius-aware pruning and tile-based rasterization for efficient rendering. To enable semantic control, each Gaussian encodes a compressed CLIP-aligned language embedding, forming a compact and queryable 3D representation. At inference time, prompt-based semantic filtering is applied directly in 3D space, without relying on image-space segmentation or view-level fusion. The selected Gaussians are then converted into dense point clouds via distribution-aware sampling and clustered to estimate fruit counts. Experimental results on real orchard data demonstrate that FruitLangGS achieves higher rendering speed, semantic flexibility, and counting accuracy compared to prior approaches, offering a new perspective for language-driven, real-time neural rendering across open-world scenarios.
Index Terms—  Radiance Fields, Gaussian Splatting, Fruit Counting, Language-guided Filtering
1 Introduction
The global transition toward smart agriculture has led to increasing demand for scalable, autonomous systems that can sense, reason, and act in dynamic outdoor environments [1]. In this context, fruit counting plays a crucial role in yield prediction, robotic harvesting, and orchard monitoring, supporting critical decision-making in precision agriculture [2]. Traditional manual methods are labor-intensive, error-prone, and lack scalability. As a result, vision-based techniques have gained traction across agricultural robotics, computer vision, and remote sensing communities [3, 4].
Early approaches relied on 2D detection from RGB images using handcrafted features or deep object detectors [5, 6], which were often hampered by occlusions, varying lighting conditions, and inconsistent viewpoints [7]. Recent works have leveraged multi-view geometry, introducing 3D-aware counting pipelines. A representative example is the use of Structure-from-Motion (SfM) to obtain camera poses and fuse 2D detections into 3D space [8], enabling more accurate fruit localization and preventing double counting. While effective, these pipelines typically require instance tracking or segmentation in every frame, limiting real-time applicability and robustness across diverse orchard scenes.
With the advent of neural scene representations, 3D reconstruction and semantic understanding have increasingly shifted toward learned volumetric representations. FruitNeRF [9], for instance, pioneered the use of semantic neural radiance fields for fruit counting, integrating RGB and segmentation masks into an implicit 3D representation. However, NeRF-based methods suffer from high computational cost and long rendering times, especially in agricultural environments where real-time feedback is essential for robotic deployment. Furthermore, semantic supervision in such pipelines typically relies on frame-level binary masks, requiring model-specific training and lacking open-vocabulary generalization.
Fig. 1: Overview of the FruitLangGS pipeline.
To address the practical demands of agricultural robotics, an effective fruit counting system must not only localize fruits accurately under occlusion and semantic ambiguity, but also provide timely feedback for downstream tasks such as robot path planning, yield estimation, and harvest coordination. In this context, fast 3D rendering facilitates real-time perception and decision-making onboard field robots, while semantic controllability enables farmers or models to flexibly query different types of fruits without retraining for each crop or scene.
We argue that a promising direction is to model the scene geometry and semantics using compact and queryable structures, avoiding the inefficiencies of implicit volumetric models or per-frame processing. To this end, we propose FruitLangGS, a real-time 3D fruit counting system that unifies efficient 3D Gaussian Splatting with language-guided semantic filtering. Our contributions are summarized as follows:
  • We propose FruitLangGS, a real-time 3D fruit counting system that reconstructs orchard-scale scenes using an adaptive Gaussian splatting pipeline. By incorporating opacity-aware pruning and load-aware tile scheduling into the rasterization process, the method achieves highly efficient rendering for large and complex fruit tree structures.
  • We design a language-conditioned semantic encoding module that embeds compressed CLIP-derived features into each 3D Gaussian. This enables open-vocabulary filtering within the 3D space, facilitating precise fruit region selection based on user-defined prompts.
  • We develop a fruit-centric counting pipeline that filters Gaussians using prompt-conditioned semantic similarity, generates color-consistent point clouds through distribution-aware sampling, and performs instance-level clustering via shape-constrained 3D aggregation.
2 Related Work
2.1 Fruit Counting in Precision Agriculture
Fruit counting is a critical task in modern precision agriculture, serving as a foundation for yield prediction, harvest planning, and crop-load management. Traditional approaches relied heavily on color features in 2D images, often assuming strong color contrast between fruits and foliage [5, 6]. However, in natural orchards, unstable lighting and background clutter often limit the reliability of color-based detection. More recently, deep learning has advanced image-based counting. Bhattarai et al. [10] proposed a weakly supervised network that regresses object counts from image-level labels, simplifying annotation requirements. Ge et al. [11] integrated detection and tracking using an improved YOLO-DeepSORT pipeline [12] for tomato yield monitoring.
To improve robustness, Barnea et al. [13] proposed a shape-based 3D fruit detection method that leverages surface geometry and symmetry rather than color. Similarly, Chu et al. [14] developed an active laser-camera system for high-precision fruit localization, demonstrating millimeter-level accuracy in challenging orchard conditions. These works highlight the increasing importance of 3D spatial reasoning in fruit analysis.
2.2 Neural Representations for 3D Scene Understanding
Neural scene representations have transformed 3D reconstruction by modeling appearance and geometry through data-driven rendering techniques. A representative example is NeRF [15], which encodes a scene as a continuous radiance field using a multilayer perceptron (MLP) and synthesizes views via differentiable ray marching. Although NeRF achieves high visual fidelity, its reliance on implicit volumetric rendering leads to high computational cost and low inference throughput. To bring NeRF into practical agricultural applications, FruitNeRF [9] extends the framework with semantic supervision, constructing fruit-aware radiance fields that allow 3D fruit counting via volumetric queries. However, the computational latency and memory demand of NeRF-style models remain a major obstacle to real-time use in unstructured orchard environments.
To address these limitations, recent works have shifted toward explicit and rasterizable representations. 3D Gaussian Splatting (3DGS) [16] replaces MLPs with a collection of spatially distributed Gaussian ellipsoids, rendered using forward splatting and tile-based alpha compositing. This approach significantly improves rendering speed while maintaining high reconstruction quality, making it well-suited for real-time robotics and interactive systems. Further improvements are achieved by AdR-Gaussian [17], which introduces adaptive radius culling to remove low-opacity splats before rasterization and load-aware tile scheduling to reduce compute imbalance. These optimizations reduce GPU workload and improve scalability for large-scale scenes.
Fig. 2: Architecture of FruitLangGS. The pipeline begins on the left with multi-view RGB images captured in real orchard scenes, which are reconstructed into 3D Gaussians using adaptive splatting. This process includes opacity-aware pruning and load-aware tile scheduling to improve rendering efficiency. In the center, each Gaussian is enriched with language-aligned semantic vectors extracted via a CLIP encoder and compressed through an autoencoder under multi-level supervision (subpart, part, whole). On the right, user-defined prompts guide cosine-based semantic filtering. The selected Gaussians are sampled into dense point clouds and clustered to estimate the final fruit count.
2.3 Language-Driven 3D Perception
The rise of vision-language models has enabled open-vocabulary scene understanding across modalities. Contrastive language–image pre-training (CLIP) [18], in particular, demonstrates that textual prompts can guide image-level classification and retrieval in a zero-shot manner. Extending this capability to 3D perception, several methods have embedded CLIP-derived features into spatial representations to enable prompt-driven querying in reconstructed scenes.
Early efforts typically relied on NeRF-based frameworks to learn 3D-consistent language fields. However, the volumetric nature of NeRF imposes high computational cost and slow inference. To address this, recent work has explored more efficient representations. LangSplat [19] proposed the first 3D language field modeled with Gaussian splatting, where the method employs a scene-specific autoencoder to reduce memory overhead and leverages the Segment Anything Model (SAM) [20] to inject hierarchical semantics at multiple levels. These innovations enable accurate and efficient open-vocabulary querying within 3D scenes while preserving clear semantic boundaries and high rendering performance.
3 Method
FruitLangGS is proposed as a framework for real-time, open-vocabulary fruit counting in orchard environments. The overall pipeline is shown in Fig. 2. We detail each module in the following sections.
3.1 Scene Reconstruction with Adaptive Splatting
In orchard-scale fruit counting, a suitable scene representation must reconcile two contrasting properties: efficient rendering and fine-grained spatial detail. The environment is composed of wide tree canopies, irregular branches, and dense fruit clusters that vary significantly in geometric scale and distribution. To model this variability while supporting high frame rates, we adopt an explicit point-based representation built on 3D Gaussian splatting.
We represent the scene as a set of $N$ anisotropic Gaussians $\{\mathcal{G}_{i}\}_{i=1}^{N}$, where each Gaussian $\mathcal{G}_{i}$ is parameterized as:
$\mathcal{G}_{i}=\left\{\boldsymbol{\mu}_{i},\Sigma_{i},\mathbf{c}_{i},\alpha_{i},\mathbf{f}_{i}\right\},$ (1)
where $\boldsymbol{\mu}_{i}\in\mathbb{R}^{3}$ denotes the 3D center position of the Gaussian, $\Sigma_{i}\in\mathbb{R}^{3\times 3}$ is a full covariance matrix encoding the Gaussian’s spatial extent and orientation, typically factorized as $\Sigma_{i}=\mathbf{R}_{i}\text{diag}(\boldsymbol{\sigma}_{i}^{2})\mathbf{R}_{i}^{\top}$ with rotation matrix $\mathbf{R}_{i}\in SO(3)$ and scale vector $\boldsymbol{\sigma}_{i}\in\mathbb{R}^{3}$, $\mathbf{c}_{i}\in\mathbb{R}^{3}$ is the RGB color, $\alpha_{i}\in[0,1]$ is the learnable opacity controlling blending behavior, and $\mathbf{f}_{i}\in\mathbb{R}^{d}$ is a learnable feature vector used in semantic embedding (see Sec. 3.2).
Given a set of calibrated RGB views $\{I_{k}\}$ with camera intrinsics and extrinsics, we optimize the parameters of $\{\mathcal{G}_{i}\}$ to minimize a multi-view photometric loss. The optimization process proceeds by projecting the Gaussians into each camera view using the projection function $\pi_{k}$, computing their influence in image space, and compositing them using alpha blending:
$\mathbf{u}_{i}^{k}=\pi_{k}(\boldsymbol{\mu}_{i}),\quad I_{k}\approx\sum_{i}\alpha_{i}\cdot\mathcal{K}(\mathbf{u}_{i}^{k}-\mathbf{p})\cdot\mathbf{c}_{i},$ (2)
where $\mathcal{K}$ is the screen-space Gaussian kernel and $\mathbf{p}$ is the target pixel.
To maintain high rendering efficiency, this module introduces two optimizations tailored to the spatial characteristics of fruit scenes. First, it applies opacity-aware pruning to reduce splatting redundancy. For each Gaussian–tile pair $(i,j)$, we define a visibility score:
$s_{ij}=\alpha_{i}\cdot\mathbb{I}[\pi(\boldsymbol{\mu}_{i})\in T_{j}],$ (3)
where $T_{j}$ is the $j$-th tile in screen space. Only Gaussians with $s_{ij}\geq\tau$ are retained for rasterization, effectively removing contributions from splats with negligible visibility. This step is implemented as a batched filtering operation using visibility masks in our training pipeline.
Second, this module adopts a tile-aware scheduling strategy to address the spatial imbalance caused by clustered fruits. To formalize this imbalance, we estimate the load $\ell_{i}$ for each pixel thread $i$ as the number of Gaussians that contribute non-trivially to its final color:
$\ell_{i}=\sum_{j\in N}\mathbb{I}(G_{j},p_{i}),$ (4)
where $\mathbb{I}(G_{j},p_{i})$ is an indicator function that equals 1 if Gaussian $G_{j}$ contributes to pixel $p_{i}$ and 0 otherwise.
Regions with high fruit density tend to yield large $\ell_{i}$, leading to rendering stalls due to sequential alpha blending within the thread. Conversely, low-density regions underutilize computational threads. To balance this, we dynamically allocate rendering bandwidth by downsampling contributions to high-load pixels and redistributing those savings to low-load tiles. This coarse-grained balancing ensures tile groups have approximately equal cumulative $\ell_{i}$, which improves scheduling efficiency and reduces latency variation across the screen. The inclusion of per-Gaussian features $\mathbf{f}_{i}$ further prepares each reconstructed point for semantic conditioning.
3.2 Language-Conditioned Semantic Embedding
Accurate fruit counting often requires distinguishing visually similar objects such as apples, leaves, and branches. Relying solely on geometry and color is insufficient due to occlusions, lighting variability, and inter-object similarity. To address this, this module enrich each 3D Gaussian with a language-conditioned feature vector that enables prompt-driven semantic selection. Our goal is to embed natural language semantics into each Gaussian such that prompt-based filtering can later isolate target instances, as detailed in Sec. 3.3.
This module begins by encoding textual semantics using CLIP. For each training image, we apply a pretrained segmentation model to obtain a set of object masks at three hierarchical levels: subpart ($s$), part ($p$), and whole ($w$). Given an input image $I_{t}$ and mask $M^{l}_{t}(v)$ centered at pixel $v$ for level $l$, the corresponding CLIP feature is extracted as:
$L^{l}_{t}(v)=V(I_{t}\odot M^{l}_{t}(v)),$ (5)
where $V(\cdot)$ denotes the CLIP encoder and $\odot$ is element-wise masking.
To reduce memory consumption and improve runtime efficiency, we compress these $512$-dimensional CLIP vectors into a lower-dimensional latent space. We train a lightweight scene-specific autoencoder $\Psi\circ E$ using the loss:
$\mathcal{L}_{\text{AE}}=\sum_{l\in\{s,p,w\}}\sum_{t=1}^{T}d_{\text{ae}}(\Psi(E(L^{l}_{t}(v))),L^{l}_{t}(v)),$ (6)
where $d_{\text{ae}}$ is a combination of $L_{1}$ and cosine distance. The encoder $E$ outputs compressed vectors $H^{l}_{t}(v)=E(L^{l}_{t}(v))\in\mathbb{R}^{d}$ with $d\ll 512$.
Each Gaussian $\mathcal{G}_{i}$ is augmented with semantic codes $f_{i}^{l}\in\mathbb{R}^{d}$ for $l\in\{s,p,w\}$, learned by minimizing the discrepancy between rendered features and projected semantic targets. Given the Gaussian rendering function for semantic features:
$F^{l}(v)=\sum_{i\in\mathcal{N}}f_{i}^{l}\alpha_{i}\prod_{j<i}(1-\alpha_{j}),$ (7)
where $\alpha_{i}$ is the opacity of Gaussian $i$ and $\mathcal{N}$ denotes the ordered Gaussians within the tile, the semantic loss is:
$\mathcal{L}_{\text{sem}}=\sum_{l\in\{s,p,w\}}\sum_{t=1}^{T}d_{\text{sem}}(F^{l}_{t}(v),H^{l}_{t}(v)),$ (8)
with $d_{\text{sem}}$ again defined as a weighted sum of $L_{1}$ and cosine loss.
This embedding process is integrated into the main training loop by extending each $\mathcal{G}_{i}$ to contain language features in addition to geometry and appearance. In practice, we set $d=3$, allowing each Gaussian to store a compact semantic vector per level. This low-dimensional encoding preserves discriminative capacity while significantly reducing storage and computation overhead during training and inference.
To jointly optimize geometry, appearance, and semantics, we define the overall objective as:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{rgb}}+\lambda_{\text{sem}}\mathcal{L}_{\text{sem}}+\lambda_{\text{reg}}\mathcal{L}_{\text{reg}},$ (9)
where $\mathcal{L}_{\text{rgb}}$ denotes the multi-view photometric reconstruction loss introduced in Sec. 3.1, $\mathcal{L}_{\text{sem}}$ is the semantic embedding loss from Eq. (6), and $\mathcal{L}_{\text{reg}}$ includes standard regularization terms over Gaussian parameters. Scalars $\lambda_{\text{sem}}$ and $\lambda_{\text{reg}}$ control the relative influence of each term.
The result is a scene composed of language-augmented Gaussians capable of open-vocabulary prompt matching. These features are optimized jointly with geometry and color, making them tightly coupled with fruit topology. Unlike prior 2D mask injection strategies, our approach maintains 3D consistency and enables direct semantic filtering at the Gaussian level without view-dependent heuristics.
3.3 Prompt-Based Filtering and Counting
Following the semantic embedding stage, each Gaussian $\mathcal{G}_{i}$ contains a compact language-aligned feature vector $\mathbf{f}_{i}\in\mathbb{R}^{d}$ that enables open-vocabulary querying. In this stage, we apply a two-step filtering process to extract fruit-relevant Gaussians based on user-specified textual prompts.
Given a set of positive prompts $\mathcal{P}^{+}=\{p_{1},\dots,p_{m}\}$ and optional negative prompts $\mathcal{P}^{-}=\{n_{1},\dots,n_{k}\}$, we first encode them using a pretrained CLIP text encoder into unit-normalized features $\mathbf{t}_{1}^{+},\dots,\mathbf{t}_{m}^{+}$ and $\mathbf{t}_{1}^{-},\dots,\mathbf{t}_{k}^{-}\in\mathbb{R}^{d}$. The cosine similarity between a Gaussian feature $\mathbf{f}_{i}$ and the prompt set is computed as:
$s_{i}^{+}=\max_{j}\cos(\mathbf{f}_{i},\mathbf{t}_{j}^{+}),\quad s_{i}^{-}=\max_{j}\cos(\mathbf{f}_{i},\mathbf{t}_{j}^{-}),$ (10)
where $\cos(\cdot,\cdot)$ denotes cosine similarity. We retain only the Gaussians that satisfy the dual-threshold filtering criterion:
=$\mathcal{M}=\left\{i\mid s_{i}^{+}>\tau_{\text{pos}}\;\land\;s_{i}^{-}<\tau_{\text{neg}}\right\},$ (11)
where $\tau_{\text{pos}}$ and $\tau_{\text{neg}}$ are manually selected thresholds. This process yields a filtered Gaussian subset that is both semantically relevant and exclusive of distractors such as leaves or branches.
This module next converts the selected Gaussians $\mathcal{G}_{i}$ for $i\in\mathcal{M}$ into a dense colored point cloud $\mathcal{P}$ for geometric clustering. Each Gaussian is sampled with a number of points proportional to its volume:
$N_{i}\propto\det(\Sigma_{i})^{\frac{1}{2}}\cdot\alpha_{i},$ (12)
where $\Sigma_{i}$ is the covariance matrix and $\alpha_{i}$ is the opacity. The sampled points inherit RGB values from $\mathbf{c}_{i}$ and normals are optionally estimated via local PCA. This strategy preserves spatial fidelity while avoiding oversampling flat or redundant regions.
To obtain fruit instances, we apply hierarchical clustering to the resulting point cloud $\mathcal{P}$. First, density-based spatial clustering of applications with noise (DBSCAN) [21] is used to obtain coarse spatial groups with parameters $\epsilon$ and min_samples adapted to fruit scale. Clusters with insufficient density are pruned. Then, we refine each cluster via template-guided splitting. For each candidate region, we compare its alpha shape volume and Hausdorff distance against a fruit template point cloud, applying iterative alignment and splitting if mismatch is detected. Let $\mathcal{C}=\{C_{1},\dots,C_{r}\}$ denote the resulting set of clusters. The fruit count is then estimated as:
$\hat{N}_{\text{fruit}}=\sum_{j=1}^{r}\gamma(C_{j}),$ (13)
where $\gamma(\cdot)$ returns the number of fruit instances fitted to cluster $C_{j}$ based on shape matching. This procedure allows robust estimation even in cases of tight occlusion, size variation, or multiple fruits in a single region.
The full pipeline from prompt to count operates directly in 3D, without requiring 2D segmentation masks, multi-view voting, or object detection heuristics. This design enables FruitLangGS to generalize across fruit species and supports prompt-level flexibility for downstream tasks.
Table 1: Semantic filtering, point sampling, and clustering configurations used across all experiments.
Component Configuration
Prompt encoder CLIP-ViT-B/16
Feature compression Autoencoder (512 $\rightarrow$ 3 dimensions per semantic level)
Similarity thresholds $\tau_{\text{pos}}=0.2255$, $\tau_{\text{neg}}=0.26125$
Sampling method Proportional to $\det(\Sigma_{i})^{1/2}\cdot\alpha_{i}$
Target points per tree $\sim$2.1M points
Sampling options Minimum opacity: 0.05; Color quality: ultra; Cleaning: enabled
Clustering algorithm DBSCAN with fruit-shape template refinement
Instance types Single, small-fruit, compound (split by Hausdorff distance + alpha shape volume)
Table 2: Reconstruction quality and rendering speed comparison across four orchard scenes. Each row reports one metric; arrows indicate whether higher ($\uparrow$) or lower ($\downarrow$) is better. Bold denotes the best result per row.
Dataset Metric NeRF [15] FruitNeRF [9] Taming3DGS [22] 3DGS [16] FruitLangGS (Ours)
Tree01 SSIM $\uparrow$ 0.401 0.414 0.429 0.539 0.517
PSNR $\uparrow$ 16.402 16.610 16.616 18.07 17.937
LPIPS $\downarrow$ 0.713 0.699 0.616 0.372 0.440
FPS $\uparrow$ $<$0.2 0.236 277.91 144.78 988.86
Tree02 SSIM $\uparrow$ 0.448 0.455 0.598 0.632 0.669
PSNR $\uparrow$ 17.907 18.028 19.783 20.708 20.176
LPIPS $\downarrow$ 0.670 0.684 0.383 0.275 0.328
FPS $\uparrow$ $<$0.2 0.285 88.87 103.59 664.44
Tree03 SSIM $\uparrow$ 0.346 0.387 0.524 0.622 0.599
PSNR $\uparrow$ 15.323 15.566 16.654 17.746 17.778
LPIPS $\downarrow$ 0.686 0.687 0.431 0.299 0.340
FPS $\uparrow$ $<$0.2 0.286 469.75 103.43 660.29
Fuji-SfM SSIM $\uparrow$ 0.603 0.599 0.633 0.693 0.716
PSNR $\uparrow$ 19.101 19.698 19.357 20.978 21.454
LPIPS $\downarrow$ 0.313 0.295 0.404 0.305 0.295
FPS $\uparrow$ $<$0.2 5.82 196.69 26.63 306.88
4 Experiments and Results
This article conducts extensive experiments to evaluate the performance of FruitLangGS in real-world fruit counting scenarios. To ensure reproducibility and allow comparison with previous work, we adopt the real-world fruit tree datasets released by FruitNeRF [9], which represent the only large-scale, open-source benchmark for per-tree 3D fruit counting.
4.1 Datasets and Implementation Details
We evaluate FruitLangGS on two real-world subsets from the FruitNeRF dataset [9]: a manually annotated 3-tree subset (Tree01–03) and the larger Fuji-SfM set with 11 additional trees. Each Tree01–03 sequence contains around 350 high-resolution posed RGB images ($4000\times 6000$), covering both sides and elevations of the canopy. Ground-truth fruit counts are available for Tree01 to Tree03 and used for instance-level evaluation. Fuji-SfM has no per-instance annotations and is used for large-scale counting benchmarks.
All experiments are conducted on a workstation with two NVIDIA RTX 4090 GPUs, an AMD 7950X CPU, and 128GB RAM. Our pipeline is implemented in PyTorch with full GPU acceleration and runs entirely offline. For each tree, we train 3D Gaussian parameters for 30,000 iterations using a learning rate of $1.6\times 10^{-3}$ with exponential decay. The input resolution is $1000\times 1500$, and the model employs $l=3$ spherical harmonics for appearance. We enable progressive densification, opacity pruning, and regularization during training, typically converging in under 10 minutes per tree.Key hyperparameters for semantic filtering, sampling, and clustering are listed in Table 1. All configurations are fixed across experiments unless otherwise stated.
Fig. 3: Qualitative results on four orchard scenes (Tree01–03 and Fuji-SfM). From top to bottom: input RGB images, manually annotated semantic labels, language-conditioned rendering, and instance-level fruit clustering results.
Table 3: Instance counting comparison. Predicted counts and recall (%) are shown for each method. “+Ours” refers to different reconstruction backbones followed by our semantic filtering and clustering pipeline. Bold values indicate the best recall per row.
Dataset NeRF [15]+Ours 3DGS [16]+Ours FruitNeRF [9] (U-Net-B) FruitLangGS (Ours) Ground Truth
Tree01 140 (78.2%) 154 (86.0%) 173 (96.6%) 173 (96.6%) 179
Tree02 88 (77.9%) 98 (86.7%) 112 (99.1%) 110 (97.3%) 113
Tree03 136 (46.7%) 194 (66.7%) 264 (90.7%) 287 (98.6%) 291
Fuji-SfM 991 (68.1%) 1284 (88.2%) 1459 (100.3%) 1443 (99.2%) 1455
Table 4: Ablation study on the effect of language-conditioned semantic embedding on fruit counting performance. Each cell reports predicted count versus GT, followed by recall percentage.
Configuration Tree01 Tree02 Tree03 Fuji-SfM
w/ Language-Conditioned Semantic Embedding 173 / 179 (96.6%) 110 / 113 (97.3%) 287 / 291 (98.6%) 1443 / 1455 (99.2%)
w/o Language-Conditioned Semantic Embedding 141 / 179 (78.8%) 93 / 113 (82.3%) 204 / 291 (70.1%) 1182 / 1455 (81.2%)
w/ Fruit Field Replacement from [9] 158 / 179 (88.3%) 101 / 113 (89.4%) 251 / 291 (86.3%) 1327 / 1455 (91.2%)
4.2 Reconstruction Quality and Efficiency
This research first assess the 3D reconstruction performance of FruitLangGS in terms of both visual quality and rendering speed. As shown in Table 2, we compare our method against four representative baselines: NeRF [15], FruitNeRF [9], Taming3DGS [22], and 3D Gaussian Splatting (3DGS) [16]. Evaluations are conducted on three annotated trees (Tree01–03) and the full Fuji-SfM dataset. The reported metrics include structural similarity (SSIM), peak signal-to-noise ratio (PSNR), learned perceptual image patch similarity (LPIPS), and frame rate (FPS).
FruitLangGS consistently achieves the highest rendering speed across all scenes. On Tree01, our method reaches 988.86 FPS, which is over three times faster than Taming3DGS (277.91 FPS) and nearly seven times faster than 3DGS (144.78 FPS). Both NeRF and FruitNeRF operate below 0.3 FPS and are unsuitable for real-time applications. On Fuji-SfM, which involves full-tree reconstruction across 11 real-world trees, FruitLangGS maintains 306.88 FPS, significantly outperforming 3DGS (26.63 FPS) and Taming3DGS (196.69 FPS). These results demonstrate that our adaptive radius-based pruning and tile-level load balancing strategies enable real-time rendering performance, even on large and structurally complex orchard scenes.
In terms of reconstruction quality, FruitLangGS achieves results that are competitive with or superior to prior work. On Fuji-SfM, our method achieves the highest SSIM (0.716) and PSNR (21.454), along with the lowest LPIPS (0.295, tied with FruitNeRF). Performance differences are more pronounced in occlusion-heavy scenes such as Tree03. In this case, NeRF and FruitNeRF achieve SSIM scores of only 0.346 and 0.387, respectively, while LPIPS remains above 0.68 for both. This indicates poor spatial alignment and perceptual quality in cluttered foliage conditions. In contrast, our method maintains stable SSIM (0.599) and lower LPIPS (0.340), suggesting improved robustness to overlapping geometry and leaf-fruit entanglement.
4.3 Instance Counting Accuracy
Figure 3 provides qualitative visualizations of our pipeline on four representative orchard scenes. In this subsection, we evaluate the effectiveness of our full pipeline in terms of counting recall, using both manually annotated trees (Tree01–03) and the full Fuji-SfM benchmark. Table 3 compares our method with three baselines: NeRF [15]+Ours, 3DGS [16]+Ours, and FruitNeRF [9] with its best-performing U-Net-Big-model (U-Net-B) segmentation backbone.
Across all scenes, FruitLangGS achieves the most consistent and accurate instance counts. On Tree03, our method correctly identifies 287 out of 291 fruits, yielding a recall of 98.6%. This is 31.9% higher than the 66.7% recall achieved by 3DGS+Ours, and 7.9% higher than FruitNeRF. On Tree01 and Tree02, our method performs on par with FruitNeRF, with 96.6% and 97.3% recall respectively. In the Fuji-SfM datasets, FruitLangGS detects 1443 out of 1455 ground-truth instances, achieving a recall of 99.2%. This matches the overall performance of FruitNeRF (1459 detections), which reports a slightly inflated recall of 100.3%, likely due to oversegmentation or false-positive clustering around ambiguous fruit boundaries. In contrast, our method maintains recall close to 100% without over-counting, demonstrating its robustness across species, geometries, and occlusion conditions. Also, while 3DGS+Ours offers moderately improved recall due to sharper geometry, it still fails to match the consistency of our adaptive Gaussian reconstruction combined with semantic filtering and clustering.
4.4 Semantic Embedding Ablation
Table 4 presents an ablation study designed to evaluate the contribution of the language-conditioned semantic embedding module. The full method, referred to as ”with Language-Conditioned Semantic Embedding,” is described in Sec. 3.2. It augments each 3D Gaussian $\mathcal{G}_{i}$ with a compact language feature vector derived from the CLIP model [18] and compressed through a lightweight autoencoder. The first ablation variant disables this semantic embedding by setting the feature level to 0, thereby removing $\mathbf{f}_{i}$ from training and inference. The second variant replaces our embedding module with the Fruit Field representation introduced in FruitNeRF [9], and adopts its original DBSCAN-based clustering strategy.
Experimental results indicate that the semantic embedding plays a critical role in accurate fruit counting. When the embedding is removed, performance degrades substantially. On Tree03, recall drops from 98.6% (287 / 291) to 70.1% (204 / 291), corresponding to an absolute reduction of 28.5%. Tree01 and Tree02 show similar trends, with recall decreasing by 17.8% and 15.0%, respectively. This confirms that semantic guidance is essential for resolving occluded or tightly clustered instances. Across the full Fuji-SfM dataset, recall falls from 99.2% to 81.2% without semantic embedding. Replacing our language-aligned embedding with the Fruit Field approach recovers part of the gap, but still underperforms in all cases. For example, Tree03 improves to 86.3% (251 / 291), which remains 12.3% lower than our method. Similar deficits of 8.3% and 7.9% persist on Tree01 and Tree02, respectively.
5 Conclusion and Future Work
We present FruitLangGS, a real-time framework for 3D fruit counting that integrates adaptive Gaussian splatting, language-conditioned semantic embedding, and prompt-driven instance estimation. Each module is designed to address specific challenges in orchard-scale perception: efficient rendering for large scenes, open-vocabulary semantic controllability, and accurate clustering under occlusion. Our reconstruction module achieves sub-second rendering by introducing spatial pruning and tile-level scheduling, enabling over 300 FPS. The semantic embedding module encodes multi-level CLIP-aligned features into each 3D Gaussian, supporting language-based filtering without per-frame supervision. The final counting stage leverages distribution-aware sampling and geometric refinement to yield accurate fruit instance estimates. Compared to the state-of-the-art FruitNeRF, FruitLangGS offers improvements in rendering speed (up to 50× faster) and memory usage, while achieving comparable or better counting accuracy. On the Fuji-SfM benchmark, our method reaches 99.2% recall, closely matching FruitNeRF’s 100.3% but without overcounting. Importantly, FruitLangGS generalizes across species and scenes without requiring model retraining for each fruit or tree. In future work, we plan to adapt this framework to embodied AI settings for agricultural simulation and task planning, and extend the semantic vocabulary to support multi-class separation and recognition of other crops and field objects.